# IPAD: Inverse Prompt for AI Detection - A Robust and Explainable LLM-Generated Text Detector

### v1

@@ Abstract
Large Language Models (LLMs) have attained human-level fluency in text generation, which complicates the distinguishing between human-written and LLM-generated texts. This increases the risk of misuse and highlights the need for reliable detectors. Yet, existing detectors exhibit poor robustness on out-of-distribution (OOD) data and attacked data, which is critical for real-world scenarios. Also, they struggle to provide explainable evidence to support their decisions, thus undermining the reliability. In light of these challenges, we propose  IPAD (Inverse Prompt for AI Detection), a novel framework consisting of a  Prompt Inverter that identifies predicted prompts that could have generated the input text, and a  Distinguisher that examines how well the input texts align with the predicted prompts. We develop and examine two versions of  Distinguishers. Empirical evaluations demonstrate that both  Distinguishers perform significantly better than the baseline methods, with version2 outperforming baselines by 9.73% on in-distribution data (F1-score) and 12.65% on OOD data (AUROC). Furthermore, a user study is conducted to illustrate that IPAD enhances the AI detection trustworthiness by allowing users to directly examine the decision-making evidence, which provides interpretable support for its state-of-the-art detection results.
IPAD: Inverse Prompt for AI Detection - A Robust and Explainable LLM-Generated Text Detector
Zheng Chen1thanks: Authors contributed equally to this work.,Yushi Feng211footnotemark: 1,Changyang He3,Yue Deng1,Hongxi Pu4,Bo Li1 1Hong Kong University of Science and Technology, Hong Kong 2The University of Hong Kong, Hong Kong 3Max Planck Institute, Germany 4University of Michigan, United States Correspondence: zchenin@connect.ust.hk
@@ 1 Introduction
Large Language Models (LLMs), characterized by their massive scale and extensive training data  Chen et al. (2024), have achieved significant advances in natural language processing (NLP)  Ouyang et al. (2022); Veselovsky et al. (2023); Wu et al. (2025). However, with the advanced capabilities of LLMs, they are subject to frequent misused in various domains, including academic fraud, the creation of deceptive material, and the generation of fabricated information Ji et al. (2023); Pagnoni et al. (2022); Mirsky et al. (2023), which underscores the critical need to distinguish between human-written text (HWT) and LLM-generated text (LGT) Pagnoni et al. (2022); Yu et al. (2025); Kirchenbauer et al. (2023).
However, due to their sophisticated functionality, LLMs pose significant challenges in the robustness of current AI detection systems Wu et al. (2025). The existing detection systems, including commercial ones, frequently misclassify texts as HWT Price and Sakellarios (2023); Walters (2023) and generate inconsistent results when analyzing the same text using different detectors Chaka (2023); Weber-Wulff et al. (2023). Studies show false positive rates reaching up to 50% and false negative rates as high as 100% in different tools Weber-Wulff et al. (2023) when dealing with out-of-distribution (OOD) datasets.
Another critical issue with the existing AI detection systems is their lack of verifiable evidence Halaweh and Refae (2024), as these tools typically provide only simple outputs like "likely written by AI" or percentage-based predictions Weber-Wulff et al. (2023). The lack of evidence prevents users from defending themselves against false accusations Chaka (2023) and hinders organizations from making judgments based solely on the detection results without convincing evidences Weber-Wulff et al. (2023). This problem is particularly troublesome not only because the low accuracy of such systems as mentioned before, but also due to the consequent inadequate response to LLM misuse, which can lead to significant societal harm Stokel-Walker and Van Noorden (2023); Porsdam Mann et al. (2023); Shevlane et al. (2023); Wu et al. (2025). These limitations highlight the pressing need for more reliable, explainable and robust detectors.
@@
In this paper, we propose IPAD (Inverse Prompt for AI Detection), a novel framework comprising two key components as shown in Figure 1: a  Prompt Inverter that reconstructs prompts from input text, and a  Distinguisher that classifies text as HWT or LGT. We consider and examine two distinct approaches for the Distinguisher: the  Prompt-Text Consistency Verifier evaluates direct alignment between predicted prompts and input text, while the  Regeneration Comparator examines contents similarity by comparing input texts with the corresponding regenerated texts. Our framework introduces a paradigm shift in AI text detection by establishing an interpretable pipeline that reveals the underlying step-by-step reasoning process, therefore it enhances both detection robustness and explainability. Through comprehensive experiments comparing these two  Distinguishers, we demonstrate their respective strengths and limitations, providing new insights into how different text characteristics affect detection performance.
Empirical evaluations demonstrate that both  Distinguishers significantly surpass baseline methods, with the  Regeneration Comparator outperforming baselines by 9.73% (F1-score) on in-distribution data and 12.65% (AUROC) on out-of-distribution (OOD) data. Additionally, the  Regeneration Comparator exhibits better performance than the  Prompt-Text Consistency Verifier on attacked data with 3.78% (F1-score), and slightly better on OOD data with 0.13% (F1-score). Furthermore, a user study indicates that IPAD enhances the AI detection experience and trustworthiness by allowing users to directly examine its decision-making evidence, which includes the predicted prompts and regenerated texts, and hence provide transparent and interpretable support for its state-of-the-art detection results. The code will be available at https://github.com/Bellafc/IPAD-Inver-Prompt-for-AI-Detection-.
@@ 2 Methodology
In this section, we illustrate our method step by step. First, we introduce the overall workflow. After that, we demonstrate the details of supervised fine-tuning (SFT) the Prompt Inverter and Distinguisher .
@@ 2.1 Workflow
IPAD consists of a  Prompt Inverter and a  Distinguisher, both fine-tuned on Microsoft’s open model Phi3-medium-128k-instruct, which together form a complete detection workflow as illustrated in Figure 1. For the Distinguisher, we develop two models and examine them in Section  3.
The Input Text ($T$) is either human-written (HWT) or LLM-generated (LGT), and it is processed by the Prompt Inverter to predict the most likely prompt that could have generated it. This Predicted Prompt ($P$) is assumed to be the input that an LLM would have used to produce the text.
$P=f_{\text{inv}}(T)$
where $f_{\text{inv}}$ stands for Prompt Inverter.
For the next step, the Predicted Prompt ($P$) is fed into an LLM (we use ChatGPT, i.e. gpt-3.5-turbo by default, and other LLMs for evaluations), to generate a corresponding Regenerated Text ($T^{\prime}$).
$T^{\prime}=f_{\text{LLM}}(P)$
After that, we consider two Distinguishers. The first one is  Prompt-Text Consistency Verifier, in which the Input Text ($T$) and the Predicted Prompt ($P$) are passed to the model.
The  Prompt-Text Consistency Verifier determines whether the Predicted Prompt ($P$) can reasonably generate the given Input Text ($T$) using an LLM. The model outputs either a  "yes" or  "no" response. If the Predicted Prompt ($P$) is likely to produce the Input Text ($T$) when fed into the LLM, the model is expected to output  "yes", indicating that the Input Text ($T$) is likely LGT. Conversely, if the Predicted Prompt ($P$) does not align well with the Input Text ($T$), the model outputs  "no", suggesting that the Input Text ($T$) is less likely to have been generated by the LLM with the Predicted Prompt ($P$), and is therefore more likely to be HWT.
$S=f_{\text{PTCV}}(T,P)$
where $f_{\text{PTCV}}$ stands for Prompt-Text Consistency Verifier in the  Distinguisher.
The second Distinguisher is  Regeneration Comparator, which considers both the Input Text ($T$) and the Regenerated Text ($T^{\prime}$).
The  Regeneration Comparator determines whether the Input Text ($T$) aligns with the Regenerated Text ($T^{\prime}$), and then outputs either a "yes" or "no" response. If the Input Text ($T$) is LGT, the model is expected to output "yes," which indicates that both the Input Text ($T$) and the Regenerated Text ($T^{\prime}$) were generated by an LLM from similar prompts. Conversely, if the Input Text ($T$) is HWT, the model is expected to output "no," which signifies that the Input Text ($T$) is meaningfully distinct from the Regenerated Text ($T^{\prime}$) and thus unlikely to have been generated by an LLM.
$S=f_{\text{RC}}(T,T^{\prime})$
where $f_{\text{RC}}$ stands for  Regeneration Comparator in the Distinguisher.
Finally, for both Distinguishers,
$\hat{Y}=\begin{cases}\text{LGT},&\text{if }S=\text{Yes}\\ \text{HWT},&\text{if }S=\text{No}\end{cases}$
where $\hat{Y}$ is the final decision of the Input Text ($T$).
2.2 Datasets
2.2.1 Prompt Inverter
The datasets used to fine-tune the Prompt Inverter include several widely adopted resources in the field. These are:
  • Instructions-2M Morris et al. (2024), a collection of 2 million user and system prompts, from which we used 30,000 prompts.
  • ShareGPT Zhang et al. (2024b), an open platform where users share ChatGPT prompts and responses, from which we used 500 samples.
  • Unnatural Instructions Zhang et al. (2024b), a dataset of diverse, creative instructions generated by OpenAI’s text-davinci-002, from which we used 500 samples.
The first three datasets aims to enhance the general querying capability of the Prompt Inverter, and are all released under the MIT license. All the samples we used are the same to the samples randomly selected in  Zhang et al. (2024a). The last dataset aims to enhance the familiarity of the Prompt Inverter with the data of the essay to detect the LLM-generated essays, and are created and examined by  Koike et al. (2024), We specifically used the LLM-generated essays and problem statements for this supervised fine-tuning (SFT).
For all 45,400 training pairs, the format is standardized as follows: Instruction: "Predict the prompt of the Input Text." Input: {LGT} or {HWT} Output: {Corresponding prompt}.
2.2.2 Distinguishers
Given that essay data are diverse, we utilize only the OUTFOX dataset Koike et al. (2024). To adapt this dataset for training our Distinguisher, we enhance it to align with the model’s requirements. The original dataset consists of 14,400 training triplets of essay problem statements, student-written essays, and LLM-generated essays. To further process the data, we apply the Prompt Inverter to both student-written and LLM-generated essays, generating corresponding Predicted Prompts. These Predicted Prompts are then used to regenerate texts via ChatGPT, i.e. gpt-3.5-turbo.
The final dataset is structured as follows:
Distinguisher version1 - Prompt-Text Consistency verifier: Instruction:"Can LLM generate text2 through the prompt text1? " Input: text 1: {Predicted Prompt}; text 2: {LGT} (or {HWT}) Output: yes (or no)
Distinguisher version2 - Regeneration Comparator: Instruction: "Text 1 is generated by an LLM. Determine whether Text 2 is also generated by an LLM with a similar prompt." Input: text 1: {Regenerated Text}; text 2: {LGT} (or {HWT}) Output: yes (or no)
Following this procedure, we construct a total of 28,800 training samples, with an equal distribution of positive and negative examples (14,400 each).
@@ 2.3 Training
The supervised fine-tuning (SFT)  Wei et al. (2022) process is performed on a dataset comprising the above-mentioned 45,400 pairs for Prompt Inverter and 28,800 pairs for both Distinguishers. We utilize Microsoft’s open model, phi3-medium-128k-instruct, and we use low-rank adaptation (LoRA) method  Hu et al. (2022) on the LLaMA-Factory framework [^1] https://huggingface.co/papers/2403.13372 Zheng et al. (2024). We train it using six A800 GPUs for 20 hours for Prompt Inverter, 7 hours for Distinguisher version1, and 4 hours for Distinguisher version2.
@@ 3 Experiments
  • $\bullet$
    Assess the robustness of IPAD (using various LLMs as generators, comparing with other detectors, and evaluating on out-of-distribution (OOD) datasets).
  • $\bullet$
    Independently analyze the necessity and effectiveness of the Prompt Inverter and the Distinguishers.
  • $\bullet$
    Explore the explainability of IPAD (through a user study and analysis of linguistic differences between prompts generated by HWT and LGT).
@@ 3.1.1 Evaluation Baselines and Metrics
The in-distribution experiments refer to the testing results presented in  Koike et al. (2024), where the data aligns with the training data used for the IPAD  Distinguishers, thereby serving as our baseline. The OOD experiments refer to the DetectRL baseline  Wu et al. (2024), which is a comprehensive benchmark consisting of academic abstracts from the arXiv Archive (covering the years 2002 to 2017) [^2] http://kaggle.com/datasets/spsayakpaul/arxiv-paper-abstracts/data, news articles from the XSum dataset  Narayan et al. (2018), creative stories from Writing Prompts  Fan et al. (2018), and social reviews from Yelp Reviews Zhang et al. (2015). It also employs three attack methods to simulate complex real-world detection scenarios, which includes the prompt attacks, paraphrase attacks, and perturbation attacks Wu et al. (2024). All the testing sets have 1,000 samples in our experiments.
The  Area Under Receiver Operating Characteristic curve (AUROC) is widely used for assessing detection method  Mitchell et al. (2023) because it considers the True Positive Rate (TPR) and False Positive Rate (FPR) across different classification thresholds. Since our models predicts binary labels, we follow the  Wilcoxon-Mann-Whitney statistic Calders and Jaroszewicz (2007), and the formula is shown in appendix  A. The  AvgRec is the average of  HumanRec and  MachineRec. In our evaluation,  HumanRec is the recall for detecting Human-written texts, and  MachineRec is the recall for detecting LLM-generated texts Li et al. (2024). The  F1 Score provides a comprehensive evaluation of detector capabilities by balancing the model’s Precision and Recall. We use  AvgRec and  F1 on in-distribution data, and we use  AUROC for OOD data to align the test benchmarks for the same dataset.

### v2

# IPAD: Inverse Prompt for AI Detection – A Reliable and Explainable LLM-Generated Essay Detector

@@ Abstract
Large Language Models (LLMs) have attained human-level fluency in text generation, which complicates the distinction between human-written and LLM-generated texts. This increases the risk of misuse and highlights the need for reliable detectors. Yet, existing detectors exhibit poor robustness on out-of-distribution (OOD) data and attacked data, which is critical for real-world scenarios. Also, they struggle to provide interpretable evidence to support their decisions, thus undermining the reliability. In light of these challenges, we propose  IPAD (Inverse Prompt for AI Detection), a novel framework consisting of a  Prompt Inverter that identifies predicted prompts that could have generated the input text, and two  Distinguishers that examine the probability that the input texts align with the predicted prompts. Empirical evaluations demonstrate that IPAD outperforms the strongest baselines by 9.05% (Average Recall) on in-distribution data, 12.93% (AUROC) on out-of-distribution (OOD) data, and 5.48% (AUROC) on attacked data. IPAD also performs robustly on structured datasets. Furthermore, an interpretability assessment is conducted to illustrate that IPAD enhances the AI detection trustworthiness by allowing users to directly examine the decision-making evidence, which provides interpretable support for its state-of-the-art detection results.
IPAD: Inverse Prompt for AI Detection – A Reliable and Explainable LLM-Generated Essay Detector
Zheng Chen1thanks: These authors contributed equally to this work. , Yushi Feng211footnotemark: 1 , Yue Deng1 , Changyang He3 , Hongxi Pu4,Bo Li1 1 Computer Science and Engineering, Hong Kong University of Science and Technology 2 School of Computing and Data Science, The University of Hong Kong 3 Max Planck Institute for Security and Privacy 4 Computer Science, The University of Michigan zchenin@connect.ust.hk, fengys@connect.hku.hk, ydengbi@connect.ust.hk, changyang.he@mpi-sp.org, hongxi@umich.edu, bli@cse.ust.hk
@@ 1 Introduction
Large Language Models (LLMs), characterized by their massive scale and extensive training data  (Chen et al., 2024), have achieved significant advances in natural language processing (NLP)  (Ouyang et al., 2022; Veselovsky et al., 2023; Wu et al., 2025). However, with the advanced capabilities of LLMs, they are subject to frequent misused in various domains, including academic fraud, the creation of deceptive material, and the generation of fabricated information (Ji et al., 2023; Pagnoni et al., 2022; Mirsky et al., 2023), which underscores the critical need to distinguish between human-written text (HWT) and LLM-generated text (LGT) (Pagnoni et al., 2022; Yu et al., 2025; Kirchenbauer et al., 2023).
However, due to their sophisticated functionality, LLMs pose significant challenges in the robustness of current AI detection systems (Wu et al., 2025). The existing detection systems, including commercial ones, frequently misclassify texts as HWT (Price and Sakellarios, 2023; Walters, 2023) and generate inconsistent results when analyzing the same text using different detectors (Chaka, 2023; Weber-Wulff et al., 2023). Studies show false positive rates reaching up to 50% and false negative rates as high as 100% in different tools (Weber-Wulff et al., 2023) when dealing with out-of-distribution (OOD) datasets.
Another critical issue with the existing AI detection systems is their lack of verifiable evidence (Halaweh and Refae, 2024), as these tools typically provide only simple outputs like "likely written by AI" or percentage-based predictions (Weber-Wulff et al., 2023). The lack of evidence prevents users from defending themselves against false accusations (Chaka, 2023) and hinders organizations from making judgments based solely on the detection results without convincing evidences (Weber-Wulff et al., 2023). This problem is particularly troublesome not only because the low accuracy of such systems as mentioned before, but also due to the consequent inadequate response to LLM misuse, which can lead to significant societal harm (Stokel-Walker and Van Noorden, 2023; Porsdam Mann et al., 2023; Shevlane et al., 2023; Wu et al., 2025). These limitations highlight the pressing need for more reliable, explainable and robust detectors.
@@
In this paper, we propose IPAD (Inverse Prompt for AI Detection), a novel and interpretable framework for detecting AI-generated text. As illustrated in Figure 1, IPAD consists of two main components: a Prompt Inverter, which reconstructs the underlying prompts from input texts, and two Distinguishers—the Prompt-Text Consistency Verifier (PTCV), which measures the alignment between the predicted prompt and input text, and the Regeneration Comparator (RC), which compares the input with the corresponding regenerated text for consistency. By explicitly modeling the reasoning path from prompt inversion to final classification, IPAD introduces a paradigm shift in AI-generated content detection, significantly enhancing both detection robustness and user interpretability.
Empirical results show that IPAD outperforms state-of-the-art baselines by 9.05% in Average Recall on in-distribution datasets, 12.93% in AUROC on out-of-distribution (OOD) datasets, and 5.48% in AUROC under adversarial attacks. IPAD also generalizes well to structured data. A user study further reveals that IPAD improves trust and usability in detection tasks by presenting concrete decision evidence, including predicted prompts and regenerated texts. Code is available [^1] https://github.com/Bellafc/IPAD-Inver-Prompt-for-AI-Detection-/.
Our contributions can be summarized as follows:
  • We introduce a novel fine-tuned inverse-prompt-based detection framework that integrates prompt reconstruction and dual consistency evaluation.
  • We achieve superior detection performance on in-distribution, OOD, adversarially attacked, and prompt-structured datasets.
  • We demonstrate through an interpretability assessment that IPAD improves human trust and interpretability in AI text detection.
@@ 2 Methodology
2.1 Preliminaries
Modules.
Our method comprises a Prompt Inverter  $f_{\text{inv}}$, and two Distinguishers, namely the Prompt-Text Consistency Verifier (PTCV) $f_{\text{PTCV}}$ and the Regeneration Comparator (RC) $f_{\text{RC}}$. Given an input text $T$, the task is to determine whether it is human-written (HWT) or generated by an LLM (LGT). We denote by $\mathcal{D}_{\text{PI}}$ the training set for $f_{\text{inv}}$, consisting of pairs $(T,P)$ where $T$ is an LLM-generated text and $P$ is its original prompt. The two distinguishers are trained using disjoint datasets: $\mathcal{D}_{\text{LGT}}$ contains LLM-generated samples and $\mathcal{D}_{\text{HWT}}$ contains human-written ones. All components are fine-tuned using Microsoft’s Phi3-medium-128k-instruct model. [^2] https://huggingface.co/microsoft/Phi-3-medium-128k-instruct.
Softmax-Based probability for Binary Classification in LLM.
To estimate the fine-tuned model’s binary classification probability (i.e., the probability of predicting “yes” or “no”), we follow the logit-based estimation approach (Yoshikawa and Okazaki, 2023). Given the model input $x$, and the output logits $z$, the model’s probability assigned to $\hat{y}$ is computed through the softmax function $\sigma$ :
$\text{Confidence}_{\text{yes}}=P(\hat{y}=\text{``yes''}\mid x)=\sigma(z)_{\text{yes}},\quad$
$\text{Confidence}_{\text{no}}=P(\hat{y}=\text{``no''}\mid x)=\sigma(z)_{\text{no}}$
Since the fine-tuned model will only output“yes” or “no”, we further calculate the probability for this binary classification as:
$\text{Probability}_{\text{yes}}=\frac{\text{Confidence}_{\text{yes}}}{\text{Confidence}_{\text{yes}}+\text{Confidence}_{\text{no}}},\quad$
$\text{Probability}_{\text{no}}=1-\text{Probability}_{\text{yes}}$
@@ 2.2 Workflow
Our framework follows a multi-stage fine-tuning pipeline with the following four steps, as illustrated in Figure 1. The details of the datasets for fine-tuning is illustrated in Appendix  A.
Step 1: Training Prompt Inverter.
We first fine-tune a model $f_{\text{inv}}$ on dataset $\mathcal{D}_{\text{PI}}$, with the data structure shown in Figure  1. For any input text $T$, $f_{\text{inv}}$ predicts the most likely prompt $P$ that could have generated it, i.e. $P=f_{\text{inv}}(T)$. The resulting Prompt Inverter is then frozen and reused in the following downstream steps.
Step 2: Training the Prompt-Text Consistency Verifier (PTCV).
Given the predicted prompt $P$ in step 1, and the input text $T\in\{\text{HWT},\text{LGT}\}$, the verifier $f_{\text{PTCV}}$ is trained to predict whether the text $T$ could plausibly be generated by an LLM using the prompt $P$. The fine-tuning datasets $\mathcal{D}_{\text{LGT}}$ and $\mathcal{D}_{\text{HWT}}$ share the same structure, with output labels "yes" for $\mathcal{D}_{\text{LGT}}$ and "no" for $\mathcal{D}_{\text{HWT}}$, as shown in the Figure  1.
After fine-tuning this module, we applied it to the validation set and computed the probability score $p_{\text{PTCV}}=f_{\text{PTCV}}(T,P)$, where the confidence value was estimated using the softmax-based method described in Section  2.1.
Step 3: Training the Regeneration Comparator (RC).
With the same predicted prompt $P$ in step 1, we use an LLM to generate a regenerated text $T^{\prime}\leftarrow LLM(P)$. By default, the LLM we use is gpt-3.5-turbo. Then, the comparator $f_{\text{RC}}$ is trained to assess whether $T$ and $T^{\prime}$ can be generated by LLM with a similar prompt. This step uses the same dataset as in Step 2, but applies a different structural formatting, as shown in Figure  1.
After fine-tuning this module, we also applied it to the validation set and computed the probability score $p_{\text{RC}}=f_{\text{RC}}(T,P)$.
Step 4: Distinguisher Merge.
To determine the final classification, we combine the two probability scores, $p_{\text{PTCV}}$ and $p_{\text{RC}}$, obtained from Step 2 and Step 3 on the validation set. Specifically, we compute a weighted ensemble as $\hat{p}=w\cdot p_{\text{PTCV}}+(1-w)\cdot p_{\text{RC}}$, and assign the prediction $\hat{Y}=\text{LGT}$ if $\hat{p}>\tau$, or $\hat{Y}=\text{HWT}$ otherwise. The weight $w\in[0,1]$ and the threshold $\tau\in[0,1]$ are treated as hyperparameters and selected via grid search on the validation set. The selected values were $w=0.45$ and $\tau=0.54$.
Inference.
We perform inference on unseen input texts $T$ by sequentially applying the trained modules. Given an input text $T$, we first use the prompt inverter $f_{\text{inv}}$ to recover the most plausible prompt $P$. The prompt is then used to regenerate a candidate text $T^{\prime}$ via the an LLM. Next, we compute two probability scores: $p_{\text{PTCV}}$, indicating whether $T$ is consistent with $P$, and $p_{\text{RC}}$, assessing the likelihood that $T$ and $T^{\prime}$ originate from the same prompt. These scores are fused into a final decision score $\hat{p}$ using the gird-searched weight $w$, and the predicted label is determined by comparing $\hat{p}$ against the threshold $\tau$. The complete algorithm can be found in Appendix  B.
The inference procedure of the IPAD framework consists of three calls through a light-weight open-sourced LLM phi-3-medium-128k-instruct. Phi-3 is a decoder-only Transformer, whithin which, the self-attention complexity per layer is $\mathcal{O}(n^{2}\cdot d)$, where $n$ is the sequence length and $d$ is the hidden dimension (Vaswani et al., 2017). The additional api call to gpt-3.5-turbo for re-generating texts introduces fixed latency but no local computation cost. Therefore, the overall computational cost is bounded by $\mathcal{O}(3\cdot L\cdot n^{2}\cdot d+\texttt{OpenAI}_{\text{api}})$, where $L=32$ is the number of layers in phi-3 (Abdin et al., 2024), which is relatively small. All three phi-3 calls can be deployed in an Nvidia V100 GPU as the minimum requirement. This demonstrates that IPAD is not computationally expensive and can be deployed with relatively modest hardware requirements.
@@ 2.3 Training
The supervised fine-tuning  (Wei et al., 2022) process is performed on a Microsoft’s open model, phi3-medium-128k-instruct, and we use low-rank adaptation (LoRA) method  (Hu et al., 2022) on the LLaMA-Factory framework [^3] https://huggingface.co/papers/2403.13372 (Zheng et al., 2024). We train it using six A800 GPUs for 20 hours for Prompt Inverter, 7 hours for PTCV, and 9 hours for RC.
@@ 3 Experiments
  • Assess the robustness of IPAD, which includes using various LLMs as generators, comparing IPAD with other detectors, and evaluating on out-of-distribution (OOD), attacked datasets, and prompt-structured datasets.
  • Independently analyze the necessity and effectiveness of the Prompt Inverter, the PTCV, and the RC.
  • Explore the user-friendliness of IPAD through an interpretability assessment.
@@ 3.1.1 Evaluation Baselines and Metrics
The in-distribution experiments refer to the testing results presented in  (Koike et al., 2024), where the data aligns with the training data used for the IPAD, thereby serving as our baseline. This baseline assesses how the RoBERTa classifiers (base and large) (Park et al., 2021), the HC3 detector (Guo et al., 2023), and the OUTFOX detector (Koike et al., 2024) perform on standard data as well as under DIPPER (Alkanhel et al., 2023) and OUTFOX attacks.
The OOD experiments refer to the DetectRL baseline  (Wu et al., 2024), which is a comprehensive benchmark, which includes four datasets: (1) academic abstracts from the arXiv Archive (covering the years 2002 to 2017) [^4] http://kaggle.com/datasets/spsayakpaul/arxiv-paper-abstracts/data, (2) news articles from the XSum dataset  (Narayan et al., 2018), (3) creative stories from Writing Prompts  (Fan et al., 2018), and (4) social reviews from Yelp Reviews (Zhang et al., 2015). It also employs three attack methods to simulate complex real-world detection scenarios, which include (1) the prompt attacks, (2) paraphrase attacks, and (3) perturbation attacks (Wu et al., 2024). DetectRL evaluates three classifiers on the OOD dataset: DetectLLM (LRR) (Su et al., 2023), Fast-DetectGPT (Bao et al., 2023), RoBERTa Classifier (Base). We included two more strong classifiers in our evaluation DetectLLM (NPR) (Su et al., 2023) and Binoculars  (Hans et al., 2024). All the testing sets have 1,000 samples in our experiments.
We further evaluate its performance on OOD datasets with structured prompts. The LongWriter dataset (Bai et al., 2025), featuring an average prompt length of 1,501 tokens, reflects IPAD’s capability to handle long-form prompts. The Code-Feedback [^5] https://huggingface.co/datasets/m-a-p/Code-Feedback and Math datasets (Hendrycks et al., 2021) contain highly structured prompts, in contrast to typical essay-like writing. We compare IPAD with baseline detectors from DetectRL to assess its relative performance under these challenging conditions.
The  Area Under Receiver Operating Characteristic curve (AUROC) is widely used for assessing detection method  (Mitchell et al., 2023). Since our models predict binary labels, we follow the  Wilcoxon-Mann-Whitney statistic (Calders and Jaroszewicz, 2007), and the formula is shown in Appendix  C. The  AvgRec is the average of  HumanRec and  MachineRec, which refers to the recall of the Human-written texts and the LLM-generated texts (Li et al., 2024).